Circular No. 7261

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

V382 VELORUM S. N. Shore, Indiana University, South Bend; H. E. Bond and R.

Downes, Space Telescope Science Institute; S. Starrfield, Arizona State University; R. D. Gehrz, University of Minnesota; J. Krautter, Landessternwarte Heidelberg; and C. E. Woodward, University of Wyoming, write: “We observed V382 Vel with the Hubble Space Telescope’s Imaging Spectrograph on June 21.6 and Aug. 29.2 UT, using the medium-resolution échelle gratings (E140M and E230M; cf. IAUC 7192). Three exposures were obtained on each date, centered at 142.5, 197.8, and 270.7 nm with resolutions of ∼ 45 000 (E140M) and 30 000 (E230M). The June spectra displayed saturated, sharp-blue-edged P-Cyg absorption on all resonance lines with terminal velocities of order 5000 km/s for Al III 186.0-nm and Si IV 140.0-nm. C IV 150.0-nm showed a profile consistent with a linear velocity gradient. Weak iron-peak absorption was still present. By the August observations, the nova had entered the emission-line phase, and iron-peak absorption and P-Cyg profiles had vanished. The following strong emission lines have been detected: N V 124.0-nm, O I 130.4-nm, C II 133.4-nm, O V 137.1-nm, Si IV 140.0-nm, N IV] 148.6-nm, C IV 155.0-nm, [Ne IV] 160.2-nm, He II 164.0-nm, O III] 166.7-nm, Si II 181.6-nm, Al III 186.0-nm, C II 232.1-nm, Al II 267.2-nm, Mg II 280.0-nm, and O III 304.0-nm. A blend at 190.0 nm can be identified as Si III] and C III], with an intensity ratio of ∼ 0.3. All emission lines show structure consistent with a flattened, ringlike shell with thickness-to-radius ratio of ∼ 0.5 and possibly asymmetric, higher-velocity, diffuse emission (HWZI ∼ 2000 km/s). Low-ionization species such as Al II, Si III], and C III] show a strong knot at −1000 km/s, at the same velocity as the strongest Balmer line knot. Most profiles show fine structure that is consistent with knots of the sort detected in V1974 Cyg. A strong interstellar spectrum is present, including Si IV and C IV, and the velocities reported on IAUC 7192 are confirmed. The integrated flux from 119.0 to 312.0 nm was 3.0 × 10^{−7} erg cm^{−2} s^{−1} on June 21.6 and 4.6 × 10^{−8} erg cm^{−2} s^{−1} on Aug. 29.2, corrected for reddening of E(B − V) = 0.2. Continued monitoring at all wavelengths is urged. FUSE and Chandra observations are currently being scheduled.”

SUPERNOVA 1999dk IN UGC 1087 CCD V magnitudes from D. Hanžl, Brno, Czech Republic: Aug. 23.980

UT, 14.32 ± 0.03; Sept. 4.952, 14.54 ± 0.05; 9.935, 14.99 ± 0.02; 12.963, 15.02 ± 0.06.

1999 September 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green